OpenStudy (anonymous):

Write a proportion for the following problem: The tax on a book costing $46 is $5. At the same tax rate, what is the tax on a book costing $72?

ganeshie8 (ganeshie8):

46/72 = 5/?

OpenStudy (anonymous):

5/46 = x/72 x= 72*5/46 = 7.83 tax: 7.83$

OpenStudy (anonymous):

46/5 = x/72 5/x = 72/46 46/5 = 72/x 72/46 = x/5 these are my answers?

ganeshie8 (ganeshie8):

46/5 = 72/x